A large pump can drain a swimming pool in 3 hours. It takes 12 hours to drain the pool using a smaller pump. How long will it take to drain the pool if both pumps are used at the same time?
Please help

Answer by richwmiller(17219)

1/a+1/b=1/x
1/3+1/12=1/x
12/36+3/36=1/x
15/36=1/x
x=36/15
x=2.4 hours working together
check
x/a+x/b=1
x/3+x/12=1
2.4/3+2.4/12=1
0.8+0.2=1
1.0=1
ok
codewt